# Layerwise complexity-matched learning yields an improved model of cortical area V2

### v2

@@ A.1.6 BarlowTwins (E2E)
Due to the fact that there is not an available pre-trained AlexNet-based version of the Barlow Twins E2E training method from Zbontar et al. (2021), we pre-train our own version based on the standard AlexNet architecture (with BatchNorm layers after each convolution/linear layer). We pool the final ‘feature’ layer such that each image is represented by a single 256-d responses vector. As was done in Zbontar et al. (2021), before the loss computation, this vector is propagated through a standard MLP projector network with 2 hidden layers. We varied different projector sizes and found the best projector setting to be one with dimensionalities: (1024, 1024, 1024) for the 2 hidden layers and output layer. We pre-train this network for 100 epochs.
@@ A.3.1 Stimulus pre-processing
Figure A.1: Example samples of each of the 15 texture families and their corresponding spectrally-matched noise images. Taken from Freeman et al. (2013).
@@ A.4 OOD and Human Behavior Benchmark
Behavior error consistency metric. This metric, derived in Geirhos et al. (2020b) measures whether there is above-chance consistency with human per-sample recognition decisions. let $c_{h,m}(s)$ be 1 if both a human observer $h$ and $m$ decide either correctly or incorrectly on a given sample s, and 0 otherwise. The observed consistency $c_{h,m}$ is the average of $c_{h,m}(s)$ over all samples. The error consistency then measures whether this observed consistency is larger than the expected consistency given two independent binomial decision makers with matched accuracy. For the details on this exact computation, see Geirhos et al. (2020b).
@@ A.5 V1 Receptive Field Comparisons
Filter Visualization. We visualize the 64 filters of each learned model in Fig. A.2. We visualize the filters in grayscale (even though the filters operate on color channels), to draw comparisons specifically between the spatial receptive field properties. By inspection, none of the models perfectly capture the nature of real V1 receptive fields; however, they all learn a set of reasonably diverse of multi-scale oriented filters. Our model and the L2-AT model seem to better capture the number of cycles within the oriented filters, whereas the supervised network receptive fields are too high-frequency. Both LCL-V1 and Supervised models, however, seem to learn more localized blob filters than the adversarially-trained model.
Figure A.2: Comparing Learned Receptive Fields. We demonstrate that our single-layer (layerwise) trained model (LCL-V1), learns a diverse set of oriented receptive fields. All 64 filters are shown for our method, adversarial robust training, and standard supervised training. Qualitatively, both our model and the adversarially-trained network learn more low-frequency filters than the standard supervised-trained network. Original filters are RGB, but we show grayscale versions here to focus on the spatial structure comparisons.
V1 Receptive Field Properties. We briefly explore some of the canonical receptive field properties highlighted in the data collected from macaque V1 neurons in Ringach (2002). For the three models, we fit Gabor receptive fields to the filters shown in Fig. A.2 by minimizing the mean squared error between a parameterized 2-d Gabor function and the model receptive field. We remove those that were not fit well by the fitting procedure.
We first measure the spatial phase of each receptive field and compare these to the distribution of spatial phase of the macacque data (Fig. A.3). None of the models perfectly match the neural data distribution, but our LCL-V1 model seems to qualitatively best capture the relative bi-modal structure around even and odd-symmetry (0 and $\pi/2$). We next compare the receptive field structure via the method in Ringach (2002), plotting the number of cycles in each dimension of 2-d Gabor function (Nx, Ny), against each other for each filter. These two parameters control receptive field shape by changing the structure of the Gabor sub-fields. We see in Fig. A.4, that while there are some outliers, most of our model receptive fields fall within a similar distribution as the macaque V1 receptive fields. Compared with the other two models, our model does a better job capturing the distribution of ’blob-like’ filters near the origin.
@@
Figure A.4: Ny vs. Nx V1 receptive field structure. Ny vs. Nx parameters for model receptive fields and macaque V1 receptive fields (obtained via fitted Gabors to each filter). Model neurons are plotted in blue and Macaque V1 data is shown in green.
@@ A.6 Additional V2 BrainScore results
Method
V2 Ceiled Explained Variance $\uparrow$
LCL-V2
$\bm{0.417\pm 0.008}$
ResNet50 L2-AT ($\epsilon=3.0$) (Madry et al., 2017)
$0.402\pm 0.007$
AlexNet L2-AT ($\epsilon=3.0$) (Madry et al., 2017)
$0.397\pm 0.007$
VOneResNet-50 L2-AT (Dapello et al., 2020)
0.380
Bag-of-Tricks Model (Riedel, 2022)
0.356
AlexNet Supervised (Krizhevsky et al., 2012)
$0.353\pm 0.009$
ResNet-50 Supervised (He et al., 2016)
$0.320\pm 0.01$
ResNext-101 ($32\times 8d$) WSL (Mahajan et al., 2018)
0.312
Table A.4: LCL-V2 outperforms all existing models on BrainScore (including those that are non-architecturally matched). We provide more detailed comparisons in explaining V2 responses given a variety of models that are either state-of-the-art overall BrainScore models (Bag-of-Tricks and ResNext-101 WSL) or prior state-of-the-art in V2 prediction (ResNet50 L2-AT). Even compared with these other architectures and methods, LCL-V2 still outperforms the prior work. For models which we could obtain the numbers, we provide standard deviation over 10 cross-validated splits.
@@ A.7 Additional Human Behavior Results
Method
IN-1K acc. $\uparrow$
OOD acc. $\uparrow$
obs. consistency $\uparrow$
error consistency $\uparrow$
LCL-V2Net
0.527
0.492 0.643 0.211
Barlow Twins (Zbontar et al., 2021)
0.459
0.451
0.607
0.166
Supervised (Krizhevsky et al., 2012) 0.590
0.443
0.597
0.165
VOneNet (Dapello et al., 2020)
0.491
0.407
0.585
0.168
L2-Robust (Madry et al., 2017)
0.399
0.391
0.573
0.176

### v3

@@ A.1.6 BarlowTwins (E2E)
Due to the fact that there is not an available pre-trained AlexNet-based version of the Barlow Twins E2E training method from Zbontar et al. (2021), we pre-train our own version based on the standard AlexNet architecture (with BatchNorm layers after each convolution/linear layer). We pool the final ‘feature’ layer such that each image is represented by a single 256-d responses vector. As implemented in Zbontar et al. (2021), before the loss computation, this vector is propagated through a standard MLP projector network with 2 hidden layers. We varied different projector sizes and found the best projector setting to be one with dimensionalities: (1024, 1024, 1024) for the 2 hidden layers and output layer. We pre-train this network for 100 epochs.
@@ A.3.1 Stimulus pre-processing
Figure A.1: Example samples of each of the 15 texture families and their corresponding spectrally-matched noise models. (Taken from Freeman et al. (2013)).
@@ A.3.2 BrainScore neural prediction
However, medians over neurons can potentially obscure the performance of any given model so we also provide an additional figure comparing the per-neuron predictions of the LCL-V2 model against the L2-AT $(\epsilon=3.0)$ and Supervised AlexNet models.
Figure A.2: Comparing per neuron predictions via the BrainScore methodology. We compare the explained variance (normalized by an estimated noise ceiling) of each of the 103 V2 neurons of our LCL-V2 model against other high-performing models. Left: LCL-V2 vs. L2-AT comparison shows that although there are a reasonable fraction of neurons which the L2-AT network predicts better ( 40%), there is a larger fraction ( 60%) which are better explained by our model. Right: LCL-V2 vs. Supervised comparison shows that for nearly all neurons (>75%), the explained variance of our model is comparable or greatly exceeds that of the Supervised network.
@@ A.3.3 Texture Modulation Analysis
As discussed above, for the regression analysis we provide explained variance scores relative to the noise ceiling (internal consistency). This is done to avoid biasing results based on predictions for neurons that are inherently unpredictable. However, for this analysis a noise ceiling on texture modulation cannot be calculated in the same way. For simplicity, we therefore calculate the texture modulations only for biological neurons which have an noise ceiling $R$ above a threshold (0.4). This excludes the bottom 10 (out of 103) neurons from the texture modulation calculation. Note, we find the results to be relatively robust to the choice of threshold.
Finally to test significance of the Spearman correlations between texture family modulation rank correlations, we use the following method outlined in Obilor & Amadi (2018):
We first transform the correlations to assume a normal distribution via Fisher’s z-transformation Fisher et al. (1921): $z_{i}=arctanh(\rho_{i})$. Given that our sample size is N=15 (15 texture families), we can calculate the standard error on the difference of correlations: $SE_{\delta_{z}}=\sqrt{\frac{2}{N-2}}$. Then, the test statistic $z_{test}=\frac{z_{1}-z_{2}}{SE_{\delta_{z}}}.$ Finally, the p-value can be calculated as $p=2*(1-CDF(|z_{test}|))$.
@@ A.4 OOD and Human Behavior Benchmark
Behavior error consistency metric. This metric, derived in Geirhos et al. (2020b) measures whether there is above-chance consistency with human per-sample recognition decisions. let $c_{h,m}(s)$ be 1 if both a human observer $h$ and $m$ decide either correctly or incorrectly on a given sample s, and 0 otherwise. The observed consistency $c_{h,m}$ is the average of $c_{h,m}(s)$ over all samples. The error consistency then measures whether this observed consistency is larger than the expected consistency given two independent binomial decision makers with matched accuracy (see Geirhos et al. (2020b) for the details on this exact computation).
@@ A.5 V1 Receptive Field Comparisons
Filter Visualization. We visualize the 64 filters of each learned model in Fig. A.3. We visualize the filters in grayscale (even though the filters operate on color channels), to draw comparisons specifically between the spatial receptive field properties. By inspection, none of the models perfectly capture the nature of real V1 receptive fields; however, they all learn a set of reasonably diverse of multi-scale oriented filters. Our model and the L2-AT model seem to better capture the number of cycles within the oriented filters, whereas the supervised network receptive fields are tend to have high frequency and narrow bandwidth. Both LCL-V1 and Supervised models, however, seem to learn more localized non-oriented filters than the adversarially-trained model.
Figure A.3: Comparing learned receptive fields. The single-layer (layerwise) trained model (LCL-V1), learns a diverse set of oriented receptive fields. All 64 filters are shown for our method, adversarial robust training, and standard supervised training. All models learn a diverse set of filters, many of them oriented. Both our model and the adversarially-trained network learn more low-frequency filters than the standard supervised-trained network. Original filters are RGB, but we show grayscale versions here to focus on the spatial structure comparisons.
V1 Receptive Field Properties. We briefly explore some of the canonical receptive field properties highlighted in the data collected from macaque V1 neurons in Ringach (2002). For the three models, we fit a linear receptive field model to the filters shown in Fig. A.3 by minimizing the mean squared error between a parameterized 2-d Gabor function (product of a sinuoisoidal grating and a Gaussian envelope) and the model receptive field. We remove those cells that were not fit well.
We first measure the spatial phase of each receptive field and compare these to the distribution of spatial phase of the macacque data (Fig. A.4). None of the models perfectly match the neural data distribution, but our LCL-V1 model seems to best capture the relative bi-modal structure around even and odd-symmetry (0 and $\pi/2$). We next compare the receptive field structure via the method in Ringach (2002), plotting the size of the receptive field envelope transverse to, and along its oscillating stripes ($(n_{x},n_{y})$, respectively), in units of the spatial period of the underlying sinusoid. We see in Fig. A.5, that while there are some outliers, our model receptive fields are distributed similarly to the V1 receptive fields measured in macaque. Compared with the other two models, our model also does a better job capturing the distribution of ’blob-like’ filters near the origin.
@@
Figure A.5: Comparison of learned oriented receptive field sizes to those measured physiologically. Each point corresponds to a single cell (model, or biological), whose receptive field was fit with a Gabor function (product of a sinusoidal grating with a Gaussian envelope). Axes represent the standard deviation of the Gaussian envelope along the directions transverse to, and along, the sinusoidal stripes, in units of the sinusoidal period (referred to as $n_{x}$ and $n_{y}$, respectively, in Ringach (2002)). Model neurons are plotted in blue and Macaque V1 neurons in green.
@@ A.6 Additional V2 BrainScore results
Method V2 Ceiled Explained Variance $\uparrow$
LCL-V2 $\bm{0.417\pm 0.008}$
ResNet50 L2-AT ($\epsilon=3.0$) (Madry et al., 2017) $0.402\pm 0.007$
AlexNet L2-AT ($\epsilon=3.0$) (Madry et al., 2017) $0.397\pm 0.007$
VOneResNet-50 L2-AT (Dapello et al., 2020) 0.380
Bag-of-Tricks Model (Riedel, 2022) 0.356
AlexNet Supervised (Krizhevsky et al., 2012) $0.353\pm 0.009$
ResNet-50 Supervised (He et al., 2016) $0.320\pm 0.01$
ResNext-101 ($32\times 8d$) WSL (Mahajan et al., 2018) 0.312
Table A.4: LCL-V2 outperforms all existing models for V2 data on BrainScore (including those that are non-architecturally matched). The table compares performance in explaining macaque V2 responses for a variety of models that are either state-of-the-art BrainScore models (Bag-of-Tricks and ResNext-101 WSL) or prior state-of-the-art in V2 prediction (ResNet50 L2-AT). For those models for which we were able to obtain the full set of explained variance values, we provide standard deviation over 10 cross-validated splits.
@@ A.7 Additional Human Behavior Results
Method IN-1K acc. $\uparrow$ OOD acc. $\uparrow$ obs. consistency $\uparrow$ error consistency $\uparrow$
LCL-V2Net 0.527 0.492 0.643 0.211
Barlow Twins (Zbontar et al., 2021) 0.459 0.451 0.607 0.166
Supervised (Krizhevsky et al., 2012) 0.590 0.443 0.597 0.165
VOneNet (Dapello et al., 2020) 0.491 0.407 0.585 0.168
L2-Robust (Madry et al., 2017) 0.399 0.391 0.573 0.176